OpenStudy (anonymous):

Will give fan and medal!!! Write x^2 - 2x - 3 = 0 in the form (x - a)^2 = b, where a and b are integers.

OpenStudy (anonymous):

(x - 4)^2 = 3 (x - 3)^2 = 2 (x - 2)^2 = 1 (x - 1)^2 = 4

OpenStudy (anonymous):

What do you think the answer is?

OpenStudy (anonymous):

(x - 3)^2 = 2 or (x - 1)^2 = 4 @tom982

OpenStudy (anonymous):

It's neither of those. For an quadratic x^2+bx+c, to write it in completed square format we say it's (x+0.5b)^2+c-0.5b^2. In your equation, b=-2 and c=-3, sub these numbers in for your answer.

OpenStudy (anonymous):

(x+0.5(-2))^2+(-3)-0.5(-2)^2 like this?

OpenStudy (anonymous):

Exactly like that :)

OpenStudy (anonymous):

so now solve it, right?

OpenStudy (anonymous):

Would it be (x - 4)^2 = 3 by any chance?

OpenStudy (anonymous):

@tom982

OpenStudy (anonymous):

@ganeshie8 @jim_thompson5910

OpenStudy (anonymous):

@tom982

OpenStudy (anonymous):

How have you gone from (x+0.5(-2))^2 to (x-4)^2?